1. Baker Industries’ net income is $21,000, its interest expense is $6,000, and its tax rate is 25%. Its notes payable equals $25,000, long-term debt equals $70,000, and common equity equals $260,000. The firm finances with only debt and common equity, so it has no preferred stock. What are the firm’s ROE and ROIC? Do not round intermediate calculations. Round your answers to two decimal places.

2.A firm has a profit margin of 5.5% and an equity multiplier of 2.8. Its sales are $70 million, and it has total assets of $42 million. What is its ROE? Do not round intermediate calculations. Round your answer to two decimal places. %
Answer #1

1]

ROE = net income / common equity

ROE = $21,000 / $260,000

ROE = 8.08%

ROIC = NOPAT / (long term debt + common equity)

NOPAT = EBIT * (1 - tax rate)

EBIT = (net income / (1 - tax rate)) + interest expense

EBIT = ($21,000 / (1 - 25%)) + $6,000 = $34,000

NOPAT = $34,000 * (1 - 25%) = $25,500

ROIC = $25,500 / ($70,000 + $260,000)

ROIC = 7.73%